OpenStudy (anonymous):

The perimeter of a rectangle is 56 inches. The length is 4 more than 3 times the width. What are the dimensions of the rectangle? What is the length of one side of the rectangle? 6 inches 12 inches 22 inches 44 inches

OpenStudy (anonymous):

Let the breadth be x , then the length is 3x+4 . Perimeter = 2*(x+3x+4) = 2*(4x+4) = 56. 28= 4x+4 , x = 6. => length is 22 inches and the breadth is 6 inches.